# DoctorAgent-RL: A Multi-Agent Collaborative Reinforcement Learning System for Multi-Turn Clinical Dialogue

### v2

@@ Abstract
Large language models (LLMs) have demonstrated excellent capabilities in the field of biomedical question answering, but their application in real-world clinical consultations still faces core challenges. Single-round consultation systems require patients to describe all symptoms upfront, leading to vague diagnosis with unclear complaints. Traditional multi-turn dialogue models, constrained by static supervised learning, lack flexibility and fail to intelligently extract key clinical information. To address these limitations, we propose DoctorAgent-RL, a reinforcement learning (RL)-based multi-agent collaborative framework that models medical consultations as a dynamic decision-making process under uncertainty. The doctor agent continuously optimizes its questioning strategy within the RL framework through multi-turn interactions with the patient agent, dynamically adjusting its information-gathering path based on comprehensive rewards from the Consultation Evaluator. This RL fine-tuning mechanism enables LLMs to autonomously develop interaction strategies aligned with clinical reasoning logic, rather than superficially imitating patterns in existing dialogue data. Notably, we constructed MTMedDialog, the first English multi-turn medical consultation dataset capable of simulating patient interactions. Experiments demonstrate that DoctorAgent-RL outperforms existing models in both multi-turn reasoning capability and final diagnostic performance. This approach shows immense practical value by reducing misdiagnosis risks in time-pressured settings, freeing clinicians for complex cases, and pioneering a strategy to optimize medical resource allocation and alleviate workforce shortages.
@@ Introduction
To address these challenges, we propose DoctorAgent-RL—a multi-agent collaborative reinforcement learning framework that reformulates clinical reasoning as a Markov Decision Process (MDP), as shown in Figure 1. Within this framework: (1) A high-fidelity patient agent based on LLMs, which generates pathologically consistent responses while mimicking the diversity of real-world communication; (2) A doctor agent initialized by cloning medical actions from real consultation records and refined through reinforcement learning (RL) to master effective questioning strategies; and (3) A consultation evaluator that provides multi-dimensional rewards based on diagnostic accuracy, patient information responsiveness, and the standardization of questions. Notably, we constructed MTMedDialog, the first English multi-turn medical consultation dataset capable of simulating realistic patient interactions. This dataset enables the training and evaluation of multi-agent systems in dynamic clinical reasoning scenarios. DoctorAgent-RL redefines the diagnostic process through reward-based strategic policy optimization: The doctor agent continuously optimizes its questioning strategy within a RL framework, obtaining immediate feedback through multi-turn interactions with the patient agent, and dynamically adjusting the information-gathering path based on the comprehensive rewards provided by the consultation evaluator. This end-to-end RL fine-tuning mechanism enables LLMs to autonomously evolve interaction strategies that align with clinical reasoning logic, rather than simply imitating surface patterns in existing dialogue data. Experiments demonstrate that DoctorAgent-RL outperforms existing models in both multi-turn reasoning capability and final diagnostic performance, demonstrating practical value in assisting clinical consultations. This work demonstrates a clear pathway for developing intelligent systems that can assist in primary consultations, thereby reducing misdiagnosis risk, freeing clinicians for complex cases, and helping to optimize healthcare resource allocation.
@@
We present MTMedDialog, the first English multi-turn medical consultation dataset capable of simulating patient interactions. The dataset comprises 8,086 training samples and 2,082 test samples derived from three Chinese benchmark datasets: IMCS21 (Chen et al. 2022), CHIP-MDCFNPC (Zhang et al. 2021), and MedDG (Liu et al. 2022). The test set covers 8 major disease categories, with detailed distribution statistics provided in Supplementary Material.
@@
$R_{\text{information}}^{t}=\begin{cases}1,&\text{if the patient agent answers normally}\\ -2,&\text{if the patient agent refuses to answer}\end{cases}$ (3)
@@
$R_{\text{compliance}}^{t}=\begin{cases}-2,&\text{Question Format Violation}\\ -5,&\text{No Diagnosis within Allowed Turns}\\ 0,&\text{otherwise (i.e., protocol is followed)}\end{cases}$ (5)
@@
Following SFT, we refine the agent’s decision-making under interaction constraints using the policy optimization algorithm detailed in Section Task Formulation. To enhance robustness and mimic real-world clinical scenarios, we introduce a Dynamic Turn Budget Training Strategy. Each training episode is assigned a random dialogue turn budget (2–10 turns). The model is explicitly reminded of the remaining turns after each interaction step, encouraging efficient information gathering. This two-stage approach ensures the agent first internalizes clinical reasoning patterns via supervised learning, then refines its strategy through interactive reward optimization. Ablation studies comparing alternative training strategies (direct SFT, direct RL, fixed-turn training) are presented in the experimental section.
@@
Table  1 presents the evaluation results of the effectiveness of patient agent in simulating real-world medical interactions. In the dimension of information control, Qwen2.5-7B demonstrates optimal compliance performance, precisely constraining output content and effectively avoiding the leakage of irrelevant information. In terms of response completeness, while DeepSeek-V3 performs best, Qwen2.5-7B still maintains near-optimal levels of key information density required for doctor-patient dialogue. Notably, all tested models achieve zero error rates in the dimension of factual conflicts, fully validating the reliability of LLMs in medical knowledge. Considering both model performance and cost-effectiveness in training, we ultimately selected Qwen2.5-7B as the implementation solution for the patient agent. For a more detailed description of the dialogue flow, please refer to Supplementary Material.
Model Information Control Response Completeness Factual Conflict
DeepSeek-V3 86.4 86.1 0.0
HuatuoGPT-o1-7B 88.1 81.7 0.0
Qwen2.5-7B-Instruct 88.8 84.4 0.0
LLaMA-3.1-8B 80.3 77.2 0.0
GLM-4-9B 72.6 70.7 0.0
Table 1: Performance comparison of different models in simulating patient agent on MTMedDialog. The best results are highlighted in bold.
@@
Experimental results in Table  2 show DoctorAgent-RL achieving a comprehensive average score of 53.9%, demonstrating significant improvements over frontier models, open-source base models, and domain-specific models. More detailed experimental results can be found in Supplementary Material. The system particularly maintains stable advantages in disease types requiring in-depth consultation, proving its superior ability to simulate doctors’ clinical consultation processes.
@@
Model DSD RSD ID GSD ND CSD ED SD Avg. Score Avg. Turns
Frontier Models
GPT-4o 49.5 50.7 48.6 46.8 47.9 52.5 46.3 41.5 49.4 3.8
DeepSeek-V3 44.8 46.6 44.2 45.0 45.1 46.3 43.7 43.0 45.2 3.3
LLaMA-3.1-70B 45.4 44.5 45.6 46.1 42.1 44.9 39.5 39.0 45.0 6.0
Open-Source Base Models
GLM-4-9B 37.5 38.3 39.7 35.2 39.1 38.3 40.0 30.5 37.5 2.1
LLaMA-3.1-8B 38.8 38.7 39.3 37.6 39.0 39.6 36.3 41.5 38.8 4.8
Mistral-7B-Instruct 40.3 41.5 39.8 41.9 44.2 39.8 36.3 35.5 40.6 3.4
Qwen2.5-7B-Instruct 43.2 42.2 40.8 43.6 39.5 45.5 35.0 40.5 42.6 6.3
Domain-Specific Models
BioMistral 28.8 28.9 29.2 23.9 29.5 25.6 37.0 25.5 28.6 9.1
UltraMedical-8B 42.8 43.8 39.9 41.7 44.7 45.4 42.2 38.5 42.9 3.5
HuatuoGPT-o1-7B 43.4 44.5 44.5 41.8 43.3 49.4 45.6 37.0 43.7 2.0
AI Hospital 46.3 47.5 44.8 46.0 46.1 44.4 45.9 43.5 46.3 3.9
DoctorAgent-RL (Ours) 54.5 52.9 55.1 52.4 50.4 57.0 51.4 48.0 53.9 8.6
Table 2: Main results by disease category on MTMedDialog dataset, showing the average of diagnostic accuracy and recommendation accuracy scores. Abbreviations: DSD (Digestive System Diseases), RSD (Respiratory System Diseases), ID (Infectious Diseases), GSD (Genitourinary System Diseases), ND (Neurological Disorders), CSD (Circulatory System Diseases), ED (Endocrine Disorders), SD (Skin Diseases). Avg. Score represents the average of diagnostic and recommendation accuracy scores across all disease categories. Avg. Turns indicates the average number of interaction turns across all disease categories. Best performing metrics are highlighted in bold. The second best results are indicated with underlines.
@@
Experimental results demonstrate the significant advantages of our proposed two-stage optimization framework (SFT + RL) in medical dialogue tasks. As shown in Table 3, DoctorAgent-RL achieves a superior average score of 53.9, outperforming all baseline models. The key advantage stems from our phased strategy: (1) SFT on doctor-patient dialogues establishes reasonable questioning baselines through behavioral cloning, followed by (2) RL optimization that enables dynamic adjustment of questioning strategies for high-value information acquisition. This combined approach improves the average diagnosis and recommendation score by 25.9% compared to the base model while enhancing proactive questioning efficiency by 36.7%. Ablation studies on three critical components further validate our design:
@@
Method DSD RSD ID GSD ND CSD ED SD Avg. Score Avg. Turns
DoctorAgent-RL (Ours) 54.5 52.9 55.1 52.4 50.4 57.0 51.4 48.0 53.9 8.6
w/o Dynamic Turn 53.5 52.4 51.4 51.4 50.9 51.8 44.1 45.0 52.7 8.5
w/o SFT 48.4 48.3 48.4 48.8 49.5 50.0 48.2 43.0 48.4 5.8
w/o RL 48.3 46.5 43.6 44.5 49.0 50.2 48.5 36.0 47.4 9.0
Qwen2.5-7B-Instruct 43.2 42.2 40.8 43.6 39.5 45.5 35.0 40.5 42.6 6.3
Table 3: Performance comparison of different fine-tuning methods for Qwen2.5-7B-Instruct on MTMedDialog. The best results are highlighted in bold.
@@ 1. Information Control Rate (0-1):
The experimental results of diagnostic accuracy are shown in Table 6. DoctorAgent-RL achieved an average diagnostic accuracy of 58.9% on the MTMedDialog dataset, showing a significant advantage over the comparison models. This result indicates that multi-turn questioning trained with RL can gather more patient information and effectively improve disease identification accuracy.
The experimental results of recommendation accuracy are shown in Table 7. DoctorAgent-RL outperformed the baseline models with an average recommendation accuracy of 48.9%. Notably, the AI Hospital model showed partial advantages in the recommendation tasks for skin diseases and endocrine diseases, possibly due to the smaller test sample sizes for these diseases, which amplified the output randomness of pre-trained models. In contrast, the multi-agent collaborative architecture of AI Hospital, through parallel interactions, more easily obtained effective solutions via probability sampling. Nevertheless, DoctorAgent-RL still maintained the best overall accuracy.
@@
This section provides a detailed explanation of the dialogue flow in DoctorAgent-RL, as illustrated in Figure  9. The framework aims to simulate a realistic diagnostic process by coordinating a structured dialogue flow between two agents: the doctor and the patient. This appendix comprehensively describes the mechanisms for dialogue initiation, progression, and termination.

### v3

@@ Abstract
Large language models (LLMs) have demonstrated excellent capabilities in the field of biomedical question answering, but their application in real-world clinical consultations still faces core challenges. Single-round consultation systems require patients to describe all symptoms upfront, leading to vague diagnosis with unclear complaints. Traditional multi-turn dialogue models, constrained by static supervised learning, lack flexibility and fail to intelligently extract key clinical information. To address these limitations, we propose DoctorAgent-RL, a reinforcement learning (RL)-based multi-agent collaborative framework that models medical consultations as a dynamic decision-making process under uncertainty. The doctor agent continuously optimizes its questioning strategy within the RL framework through multi-turn interactions with the patient agent, dynamically adjusting its information-gathering path based on comprehensive rewards from the Consultation Evaluator. This RL fine-tuning mechanism enables LLMs to autonomously develop interaction strategies aligned with clinical reasoning logic, rather than superficially imitating patterns in existing dialogue data. Notably, we constructed MTMedDialog, a novel English multi-turn medical consultation dataset designed for dynamic, interactive simulations. It features comprehensive case profiles that allow a patient agent to reveal symptoms progressively based on the doctor agent’s inquiries, thus providing a more realistic training and evaluation environment compared to existing static dialogue datasets. Experiments demonstrate that DoctorAgent-RL outperforms existing models in both multi-turn reasoning capability and final diagnostic performance. This approach shows immense practical value by reducing misdiagnosis risks in time-pressured settings, freeing clinicians for complex cases, and pioneering a strategy to optimize medical resource allocation and alleviate workforce shortages.
@@ Introduction
To address these challenges, we propose DoctorAgent-RL—a multi-agent collaborative reinforcement learning framework that reformulates clinical reasoning as a Markov Decision Process (MDP), as shown in Figure 1. Within this framework: (1) A high-fidelity patient agent based on LLMs, which generates pathologically consistent responses while mimicking the diversity of real-world communication; (2) A doctor agent initialized by cloning medical actions from real consultation records and refined through reinforcement learning (RL) to master effective questioning strategies; and (3) A consultation evaluator that provides multi-dimensional rewards based on diagnostic accuracy, patient information responsiveness, and the standardization of questions. While existing datasets provide valuable examples of multi-turn medical conversations, they consist of static transcripts. This limits their use for training agents that need to learn adaptive questioning strategies through interactive simulation. To address this gap, we constructed MTMedDialog, the first English dataset specifically designed to support dynamic, stateful patient simulations. Each dialogue is grounded in a comprehensive patient profile, allowing for realistic, turn-by-turn information revelation based on the clinician agent’s diagnostic reasoning process. DoctorAgent-RL redefines the diagnostic process through reward-based strategic policy optimization: The doctor agent continuously optimizes its questioning strategy within a RL framework, obtaining immediate feedback through multi-turn interactions with the patient agent, and dynamically adjusting the information-gathering path based on the comprehensive rewards provided by the consultation evaluator. This end-to-end RL fine-tuning mechanism enables LLMs to autonomously evolve interaction strategies that align with clinical reasoning logic, rather than simply imitating surface patterns in existing dialogue data. Experiments demonstrate that DoctorAgent-RL outperforms existing models in both multi-turn reasoning capability and final diagnostic performance, demonstrating practical value in assisting clinical consultations. This work demonstrates a clear pathway for developing intelligent systems that can assist in primary consultations, thereby reducing misdiagnosis risk, freeing clinicians for complex cases, and helping to optimize healthcare resource allocation.
@@
We present MTMedDialog, a novel benchmark dataset specifically designed for training and evaluating AI agents in a simulated, multi-turn medical consultation setting. A key distinction from previous datasets is that each dialogue in MTMedDialog is grounded in a comprehensive, hidden patient profile. This structure enables a patient simulator to dynamically reveal relevant symptoms and medical history in response to a doctor agent’s questions, rather than following a fixed script. This design facilitates the development and evaluation of agents with advanced, adaptive inquiry strategies. The dataset comprises 8,086 training samples and 2,082 test samples derived from three Chinese benchmark datasets: IMCS21 (Chen et al. 2022), CHIP-MDCFNPC (Zhang et al. 2021), and MedDG (Liu et al. 2022). The test set covers 8 major disease categories, with detailed distribution statistics provided in Supplementary Material.
@@
[…]
@@
[…]
@@
Following SFT, we refine the agent’s decision-making under interaction constraints using the policy optimization algorithm detailed in the Task Formulation section. To enhance robustness and mimic real-world clinical scenarios, we introduce a Dynamic Turn Budget Training Strategy. Each training episode is assigned a random dialogue turn budget (2–10 turns). The model is explicitly reminded of the remaining turns after each interaction step, encouraging efficient information gathering. This two-stage approach ensures the agent first internalizes clinical reasoning patterns via supervised learning, then refines its strategy through interactive reward optimization. Ablation studies comparing alternative training strategies (direct SFT, direct RL, fixed-turn training) are presented in the experimental section.
@@
Figure  3 presents the evaluation results of the effectiveness of patient agent in simulating real-world medical interactions. In the dimension of information control, Qwen2.5-7B demonstrates optimal compliance performance, precisely constraining output content and effectively avoiding the leakage of irrelevant information. In terms of response completeness, while DeepSeek-V3 performs best, Qwen2.5-7B still maintains near-optimal levels of key information density required for doctor-patient dialogue. Notably, all tested models achieve zero error rates in the dimension of factual conflicts, fully validating the reliability of LLMs in medical knowledge. Considering both model performance and cost-effectiveness in training, we ultimately selected Qwen2.5-7B as the implementation solution for the patient agent. For a more detailed description of the dialogue flow, please refer to Supplementary Material.
Figure 3: Performance comparison of different models in simulating patient agent on MTMedDialog.
@@
Experimental results in Figure 4(a) show DoctorAgent-RL achieving a comprehensive average score of 53.9%, demonstrating significant improvements over frontier models, open-source base models, and domain-specific models. More detailed experimental results can be found in Supplementary Material. The system particularly maintains stable advantages in disease types requiring in-depth consultation, proving its superior ability to simulate doctors’ clinical consultation processes.
@@
Figure 4: (a) Main results by disease category on MTMedDialog dataset, showing the average of diagnostic accuracy and recommendation accuracy scores. Dark blue boxes represent Frontier Models, light blue boxes represent Open-Source Base Models, and dark gray boxes represent Domain-Specific Models. (b) Performance comparison of different fine-tuning methods for Qwen2.5-7B-Instruct on MTMedDialog. Avg. Turns indicates the average number of interaction turns across all disease categories.
Figure 5: Performance of DoctorAgent-RL in real-world interactive scenarios with actual patients.
Traditional automated evaluations and testing based on LLM-simulated patient datasets struggle to fully capture a model’s actual performance in real, dynamic conversational interactions. To comprehensively assess the end-to-end capabilities of DoctorAgent-RL in real clinical scenarios, we designed and conducted a real-world human-machine interaction evaluation. For this purpose, we recruited 20 real patients, ranging in age from their 20s to 80s, to engage in real-time online consultations with our model regarding their actual health conditions, which covered 15 common disease types. Subsequently, professional evaluators scored the complete dialogue process across multiple dimensions.
As detailed in Figure 7, DoctorAgent-RL demonstrated exceptional performance in these real-world interaction tests. The model achieved consistently high scores for both diagnostic accuracy and recommendation rationality, affirming its capacity to make reliable clinical judgments based on dynamically gathered information. Furthermore, the high scores in question quality, coupled with an efficient number of average dialogue turns, indicate that the model’s path to achieving high accuracy is both clinically logical and effective.
To further scrutinize the model’s clinical precision, we calculated the Exact Diagnostic Match Rate(ED), which measures the percentage of diagnoses that perfectly matched the ground truth. DoctorAgent-RL achieved a rate of 70%, signifying that in the vast majority of interactions, it not only provided a reasonable assessment but correctly identified the patient’s specific condition. This high rate of exact matches, combined with the strong overall performance scores, confirms that DoctorAgent-RL possesses the ability to stably, efficiently, and accurately complete clinical dialogue tasks in real-world interactive scenarios.
In-depth Analysis of Generalization and General-Purpose Capabilities
Figure 6: Qualitative comparison of dialogue outputs from DoctorAgent-RL (left column) and HuatuoGPT-o1-7B (right column) during interactions with a human user. The top row showcases performance on a specialized medical task, while the bottom row demonstrates capabilities on an open-domain general task. This comparison highlights the differences in dialogue quality, fluency, and the ability to maintain general-purpose skills after domain-specific training. In the dialogues, blue boxes indicate the human user’s input, and green boxes indicate the LLM’s output.
Having validated the model’s overall performance in real-world scenarios, we proceeded with an in-depth analysis focusing on two dimensions that are critical for evaluating the maturity of a specialized LLM: its generalization ability and its general-purpose capabilities. The former assesses the model’s capacity to handle unseen clinical cases, while the latter evaluates whether its foundational conversational skills remain intact after domain-specific training.
We began by analyzing the performance discrepancies between disease types the DoctorAgent-RL had seen versus those it had not been exposed to during its training phase. As illustrated in Table LABEL:tab:real_world_eval, the results for both groups of cases showed negligible differences across core metrics. This suggests that the model has acquired a transferable clinical reasoning framework rather than relying on rote memorization of disease-specific patterns, thereby demonstrating strong generalization capabilities.
To further evaluate the robustness and general-purpose capabilities of DoctorAgent-RL, we conducted a qualitative analysis where human evaluators interacted with multiple models across a range of scenarios, covering specialized medical inquiries, open-domain tasks (e.g., travel planning), and role-playing exercises. In our tests, we observed significant behavioral differences across different training paradigms. For instance, domain-specific SFT LLMs like BioMistral exhibited severe instruction-following deficiencies, leading to unsatisfactory performance in both medical and general tasks; its effectiveness is likely confined to a narrow distribution highly similar to its training data. In contrast, a domain-specific RL LLM like HuatuoGPT-o1-7B, despite performing well on the medical task, showed clear signs of being ”over-tuned,” resulting in rigid, template-like outputs and a severe degradation of its general-purpose conversational abilities. In contrast, DoctorAgent-RL successfully avoids this issue. It not only excels in specialized medical dialogues but also fully preserves the powerful general-purpose conversational skills of its base model, remaining fluent and efficient in non-medical tasks. The model is also robust in terms of role consistency; when instructed to switch roles and act as a ”patient,” it can describe symptoms naturally without any confusion from its core ”doctor” training. Figure 6 shows a set of representative dialogues that vividly illustrates the behavioral differences between our model and HuatuoGPT-o1-7B on both a medical and a travel planning task. Based on this qualitative evaluation, we conclude that DoctorAgent-RL is an intelligent agent that achieves deep domain specialization while fully preserving its general-purpose capabilities.
Human Evaluation and In-depth Analysis of Multi-Model Approaches on MTMedDialog
Figure 7: Human randomly selected 100 data samples to compare the performance of different models on the MTMedDialog dataset.
To conduct a more detailed evaluation of the comprehensive performance of different models in simulating real-world clinical consultation scenarios, we performed a human evaluation. Specifically, we randomly selected 100 samples from the MTMedDialog test set and then anonymized and shuffled the dialogues generated by all models for these samples. Evaluators, blinded to the model origins, independently scored the dialogues against the ground truth references. The scores were then mapped back to the respective models using their IDs to ensure the objectivity and fairness of the evaluation. Evaluation dimensions included Diagnostic Accuracy (HR-D), Recommendation Reasonableness (HR-R), Question Quality (HR-Q), and Information Coverage (HR-IC). The detailed results are presented in Table LABEL:tab:HR.
The human evaluation results clearly indicate that our proposed DoctorAgent-RL outperforms all other models, achieving the highest scores in the three core metrics of Diagnostic Accuracy, Question Quality, and Information Coverage. It comprehensively surpassed frontier closed-source models, open-source base models, and other domain-specific models. This demonstrates that our method not only excels in the final diagnostic task but also holds a significant advantage in the strategic effectiveness of the consultation process.
In contrast, other models generally showed significant deficiencies in instruction following and interaction quality. Frontier models, representing the state-of-the-art in large language models, demonstrated strong overall capabilities. Among them, GPT-4o achieved high scores across the board, making it the best alternative to our model. DeepSeek-V3 excelled in question quality but sometimes violated the one-question-per-turn instruction. LLaMA-3.1-70B showed stable performance but lagged slightly behind the other two, similarly posing multiple questions in a single turn. While this increased information coverage, it also highlighted its weakness in instruction following. Open-source base models commonly suffered from poor instruction adherence and limited medical knowledge. For instance, Qwen2.5-7B-Instruct tended to provide a diagnosis directly rather than gathering information through questions and repeatedly violated the single-question rule. Mistral-7B-Instruct had the worst performance in instruction following, often failing to ask questions or producing repetitive, irrelevant output. Although LLaMA-3.1-8B could formulate relatively professional questions, its diagnostic accuracy was low due to insufficient internal knowledge. The advantages of DoctorAgent-RL were particularly prominent among domain-specific models. Other models in this category exhibited numerous issues. HuatuoGPT-o1-7B and BioMistral-7B displayed severe problems with instruction following, even exhibiting ”role reversal” by mimicking patient symptom descriptions. HuatuoGPT-o1-7B rarely asked questions and often included its ”overthinking process” in the answers. BioMistral-7B had the highest average number of turns, but this was due to incessant, repetitive, and irrelevant questioning, leading to inefficiency and the lowest scores across all metrics. This phenomenon strongly proves that the key to effective multi-turn dialogue is question quality, not quantity. UltraMedical-8B severely violated the single-question instruction by asking five to six questions at once, resulting in too few turns for in-depth, progressive information gathering. AI Hospital, a relatively high-performing domain-specific model, achieved performance close to that of frontier models but still showed a clear gap compared to our model.
In summary, our DoctorAgent-RL model, through SFT and RL, has learned to conduct high-quality, multi-turn interactions while adhering to the ”one question at a time” instruction. The model is able to act like a real doctor, effectively gathering key information through a progressive and logical chain of questions, thereby achieving the highest diagnostic accuracy and information coverage.
@@
Experimental results demonstrate the significant advantages of our proposed two-stage optimization framework (SFT + RL) in medical dialogue tasks. As shown in Figure 4(b), DoctorAgent-RL achieves a superior average score of 53.9, outperforming all baseline models. The key advantage stems from our phased strategy: (1) SFT on doctor-patient dialogues establishes reasonable questioning baselines through behavioral cloning, followed by (2) RL optimization that enables dynamic adjustment of questioning strategies for high-value information acquisition. This combined approach improves the average diagnosis and recommendation score by 25.9% compared to the base model while enhancing proactive questioning efficiency by 36.7%. Ablation studies on three critical components further validate our design:
@@ 1. Information Control Rate (0-1):
The experimental results of diagnostic accuracy are shown in Table Detailed Comparative Performance Evaluation of Multi-Model Approaches on MTMedDialog. DoctorAgent-RL achieved an average diagnostic accuracy of 58.9% on the MTMedDialog dataset, showing a significant advantage over the comparison models. This result indicates that multi-turn questioning trained with RL can gather more patient information and effectively improve disease identification accuracy.
The experimental results of recommendation accuracy are shown in Table Detailed Comparative Performance Evaluation of Multi-Model Approaches on MTMedDialog. DoctorAgent-RL outperformed the baseline models with an average recommendation accuracy of 48.9%. Notably, the AI Hospital model showed partial advantages in the recommendation tasks for skin diseases and endocrine diseases, possibly due to the smaller test sample sizes for these diseases, which amplified the output randomness of pre-trained models. In contrast, the multi-agent collaborative architecture of AI Hospital, through parallel interactions, more easily obtained effective solutions via probability sampling. Nevertheless, DoctorAgent-RL still maintained the best overall accuracy.
@@
This section provides a detailed explanation of the dialogue flow in DoctorAgent-RL, as illustrated in Figure  14. The framework aims to simulate a realistic diagnostic process by coordinating a structured dialogue flow between two agents: the doctor and the patient. This appendix comprehensively describes the mechanisms for dialogue initiation, progression, and termination.